OpenStudy (anonymous):

Apples in a french market cost 2.5 Euros per kilogram. what is the price in dollars per pound. (Based on $ 1.256 - Euros .7965)

OpenStudy (anonymous):

\[\frac{ 2.5 euros }{ 1 kg } \times \frac{ ? U.S $ }{ ? euro } \times \frac{ 1 kg }{ ? lb }\]

OpenStudy (anonymous):

and, 1 kg=2.20462 lb

OpenStudy (anonymous):

OK I did this: 2.5/1 * $1.256/.7965 E * 1KG/2.204 =3.14/178= 1.76. I don't this this is correct.

OpenStudy (anonymous):